is anyone here good at math !?!? I NEED HELP :)
Let
p(x) = x^3-3x^2-2x+1.
p(-2) can be computed via synthetic division by filling in for the letters a through g in the table

1 -3 -2 1
b d f
a c e g




where
=a,
=b,
=c,
=d,
=e,
=f,
=g.
Moreover, p(-2) = and p(x) can be written as
p(x) = (x+2)q(x) + r
where
q(x) = is a quadratic polynomial and
r= is a real number.

Answer by stanbon(75887)

Let
p(x) = x^3-3x^2-2x+1.
p(-2) can be computed via synthetic division by filling in for the letters a through g in the table
-2)....1....-3....-2....1
........1....-2...10...-16
______________________________
.........1....-5...8...|..-15

=============================================


Moreover, p(-2) = and p(x) can be written as
p(x) = (x+2)q(x) + r
where
q(x) = x^2 -5x + 8
r= -15
===================
Cheers,
Stan H.
===================